OpenStudy (anonymous):

Find the value of x. log 2 x = 3

OpenStudy (phebe):

ima help

OpenStudy (anonymous):

Okay :D

OpenStudy (anonymous):

\[\log_2(x) = 3\]?

OpenStudy (phebe):

wats the prob.

OpenStudy (anonymous):

\[\large \log_b(x)=y\iff b^y=x\]

OpenStudy (anonymous):

I need help solving the problem. The choices are 5, 6,8, or 9. I'm getting a totally different answer.

OpenStudy (phebe):



OpenStudy (anonymous):

O.o is that a 6?

OpenStudy (phebe):

yes

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

\[\large \log_b(x)=y\iff b^y=x\] \[\large \log_2(x)=3\iff 2^3=x\][

OpenStudy (phebe):

its rong rit

OpenStudy (anonymous):

O.o is it 6 or 8?

OpenStudy (phebe):

8

OpenStudy (phebe):

6 wrong

OpenStudy (anonymous):

Okay, 8